OpenStudy (anonymous):

Factor 6x2 – 2x – 4. please I really need help

OpenStudy (mertsj):

1. factor out 2

OpenStudy (anonymous):

2(3x^2 - x -2)

OpenStudy (dwelsher):

then you have the two remaining factor(3x+2)(x-1) Then you have to check your work which I forget quite often

OpenStudy (whpalmer4):

\[2(3x^2-x-2)\]When you have a polynomial to factor where the highest power term has a coefficient other than 1, factoring by grouping is often effective. Multiply the coefficient of the highest power term and the coefficient of the constant term. Now find all the pairs of factors of that product. Look for a pair which adds up to the middle term's coefficient. We have 3*(-2) = -6 Factors of -6 are: -1,6 -2, 3 -3, 2 -6,1 1,-6 2,-3 3,-2 6,-1 2 and -3 add to -1, which is the coefficient of the middle term. Now we rewrite our polynomial with those two coefficients both multiplying \(x\): \[(3x^2-3x)+(2x-2)\]Factor each group: \[3x(x-1)+2(x-1)\]Notice that \((x-1)\) is a common factor of both product terms, and factor it out:\[(x-1)(3x+2)\]Test the factoring:\[(x-1)(3x+2) = 3x^2+2x-3x-2 = 3x^2-x-2\] Combine that with the \(3\) we factored out previously to get the full result: \[6x^2-2x-4 = 3(x-1)(3x+2)\]

OpenStudy (whpalmer4):

Sorry, combine that with the 2 we factored out previously to get the full result: \[6x^2-2x-4 = 2(x-1)(3x+2)\]